OpenStudy (anonymous):

The sum of three numbers is 50. The second number is three times the first number, and the third number is twice the second number. What are the numbers?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

let numbers be a, b, and c the sum of three numbers is 50 a+b+c=50

pooja195 (pooja195):

the second number b is three times the first number a, so you have b=3a

pooja195 (pooja195):

and the third c number is twice the second b, so you have c=2b=2(3a)=6x 3a+a+6a=50 solve for a

OpenStudy (anonymous):

A=5

pooja195 (pooja195):

good so one of the numbers is 5

pooja195 (pooja195):

now find b We have b=3a so we know a=5 so b=3(5) what is 3 x 5 ?

OpenStudy (anonymous):

15

pooja195 (pooja195):

Ok so now we have 5,15 we need to find c c=6a (as i explained above) c=6(5) 6 x 5 =?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

A=5 B=15 C=30

pooja195 (pooja195):

Good so our answer would be 5,15,30 to check if we did it correctly we can retriceall the numbers and they = 50 :) Thats all it is

OpenStudy (anonymous):

thank you here is your medal

pooja195 (pooja195):

:] thanks!

OpenStudy (anonymous):

i hope you can help me again soon

pooja195 (pooja195):

Im sure i can :)

OpenStudy (anonymous):

bye